OpenStudy (anonymous):

f(x)=1/x and g(x)=x^2-4x what two numbers are not in the domainof f o g?

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

f o g = 1/ (x^2 - 4x)

OpenStudy (anonymous):

thx bro @welshfella

OpenStudy (welshfella):

= 1 / (x(x - 4))

OpenStudy (anonymous):

do i have to simplify that at all?

OpenStudy (welshfella):

no what you need to do now is to work out what 2 values of x will make the denominator = zero. These 2 numbers will not be in the domain of f o g

OpenStudy (anonymous):

(0,4)

OpenStudy (welshfella):

basically you need to solve x(x - 4) = 0 yea you got it

OpenStudy (welshfella):

good gtg

OpenStudy (anonymous):

ok ttyl